OpenStudy (anonymous):

what is the degree of the monomial? ^7x^8 7 8 15 56

OpenStudy (anonymous):

arent you missing something there?, the starting ^7 makes no sense alone...

OpenStudy (anonymous):

thats what i thought but no thats how the answer there is shown. maybe its 7x8 ? I'm not really sure.

OpenStudy (anonymous):

mmm.. anyway, the degree is the value of the exponent of the x power... so in this case we have "x to the _____ power"

OpenStudy (anonymous):

so snice there both x i believe i should add them? which is 15?

OpenStudy (anonymous):

but there is only one x, right? i read ^7x^8, which reads as "something missing to the 7th power times x to the eigth power"

OpenStudy (anonymous):

yeah theres only one x :o so does that mean the x only belongs to 8?

OpenStudy (anonymous):

yeap, x is powered to 8, so the degree should be 8 :)

OpenStudy (anonymous):

thank you so much! c: can you help me with a few more questions?

OpenStudy (anonymous):

lets see if i can... shoot! :D

OpenStudy (anonymous):

Solve the equation by completing the square.Round to the nearest hundredth if necessary. x^2+2x=8 x=2,4 x=-2,4 x=2,-4 x=-2,-4

OpenStudy (anonymous):

ok! do you recall what completing squares is?

OpenStudy (anonymous):

I don't believe I do

OpenStudy (anonymous):

and I have a question about the first question I asked so i multiply 7*8? right?

OpenStudy (anonymous):

it consists in taking the equation to something like this: \[(x-a)^2 = b\] so it is easier to solve

OpenStudy (anonymous):

in the first one, it is not clear what the 7 is there for... if we assume it is not in the power of the x, then we dont need it for the answer

OpenStudy (anonymous):

oh okay thank you!

OpenStudy (anonymous):

and for the 2nd question is the answer a?:o

OpenStudy (anonymous):

i found it is the third one... 2,-4

OpenStudy (anonymous):

thank you! c: i have a few more is that alright?